# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

### v2

@@ Abstract
Signed networks—where edges are labeled as positive or negative to represent friendly or antagonistic interactions—offer a natural framework for analyzing polarization, trust, and conflict in social systems. Detecting meaningful group structures in such networks is crucial for understanding online discourse, political divisions, and trust dynamics. A key challenge is to identify communities that are internally cohesive and externally antagonistic, while allowing for neutral or unaligned vertices. In this paper, we propose a method for identifying $k$ polarized communities that addresses a major limitation of prior methods: their tendency to produce highly size-imbalanced solutions. We introduce a novel optimization objective that avoids such imbalance. In addition, it is well known that approximation algorithms based on local search are highly effective for clustering signed networks when neutral vertices are not allowed. We build on this idea and design the first local search algorithm that extends to the setting with neutral vertices while scaling to large networks. By connecting our approach to block-coordinate Frank-Wolfe optimization, we prove a linear convergence rate, enabled by the structure of our objective. Experiments on real-world and synthetic datasets demonstrate that our method consistently outperforms state-of-the-art baselines in solution quality, while remaining competitive in computational efficiency.
@@ 1 Introduction
The main contributions of this paper are as follows: (i) We propose a novel formulation for the PCD problem that encourages more balanced communities, addressing a key limitation of previous work that typically optimize polarity (DBLP:conf/nips/TzengOG20, ) (see Eq. 2). As demonstrated in our experiments, optimizing polarity often leads to clustering solutions with multiple empty clusters. The importance of promoting balanced communities (to avoid trivial solutions where all objects are placed in one or few clusters) is well established in the graph clustering literature DBLP:conf/cikm/ChiangWD12 . In Appendix E, we expand on this, and provide some examples of practical scenarios where (reasonably) balanced communities are favorable in our context. We note that DBLP:journals/ml/GulloMT24 also proposes an objective for PCD called $\gamma$-polarity aimed at addressing cluster imbalance; however, it is restricted to the case of $k=2$ clusters. In contrast, our objective supports an arbitrary number of clusters $k$. Nonetheless, we compare to their method experimentally for $k=2$, and explain how it differs conceptually in Appendix F. (ii) Motivated by the effectiveness of local search-based approximation algorithms for CC (and many other machine learning models), we propose the first scalable local search algorithm for PCD, which explicitly allows for neutral objects. (iii) We establish a linear convergence rate of our local search algorithm by connecting it to block-coordinate Frank-Wolfe optimization (fw, ; DBLP:conf/icml/Lacoste-JulienJSP13, ). This connection is made possible utilizing the specific structure of our proposed optimization objective. (iv) We propose techniques that allow the local search method to scale to large networks. (v) Finally, through extensive experiments on commonly used real-world and synthetic datasets in previous work on PCD, we show that our approach consistently outperforms state-of-the-art baselines in terms of (a) recovering ground-truth solutions and (b) finding high quality solutions of reasonable cluster size balance.
@@ 2.1 Correlation Clustering
We begin by noting that for CC, unlike PCD to be discussed in the next subsection, a clustering $S_{[k]}$ is a partition of $V$, meaning $V=\bigcup_{m\in[k]}S_{m}$ and each $S_{m}$ is disjoint. A notable feature of CC is its ability to automatically determine the number of clusters DBLP:conf/kdd/BonchiGL14 , but here we we focus on the $k$-constrained variant of CC (DBLP:journals/toc/GiotisG06, ) as it is most relevant to our problem. The $k$-CC problem can be defined as shown below.
Problem 1 ($k$-CC).
@@
In other words, the goal is to find a clustering that (i) maximizes intra-cluster similarities and (ii) minimizes inter-cluster similarities. In the CC literature, it is known that maximizing certain subsets of terms in Eq. 1—such as the total number of agreements, $N^{+}_{\text{intra}}+N^{-}_{\text{inter}}$—is equivalent to maximizing the full objective ethz-a-010077098 . However, this equivalence does not hold for PCD when neutral objects are allowed, as each of the four terms in Eq. 1 contributes uniquely to the decision of whether an object should be clustered or left neutral. We formally show this in Appendix A, and thus focus on the full objective in the next section when we introduce PCD. CC is known to be NP-hard DBLP:journals/ml/BansalBC04 , and many approximation algorithms have been proposed (DBLP:journals/ml/BansalBC04, ; CharikarGW05, ; DemaineEFI06, ; AilonCN08, ), with local search methods standing out for their strong performance in both clustering quality and computational efficiency (DBLP:conf/aaai/ThielCD19, ; Chehreghani22_shift, ; DBLP:journals/tmlr/AronssonC24, ).
@@ 2.2 Polarized Community Discovery
For PCD, we introduce a neutral set $S_{0}$. As a result, each object in $V$ is either assigned to one of the non-neutral clusters $S_{1},\dots,S_{k}$ or designated as neutral by placing it in $S_{0}$. Consequently, a clustering $S_{[k]}$ is no longer a partition of $V$ and we have $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ (although all clusters are still disjoint). Given this, the goal of PCD is to identify non-neutral clusters $S_{[k]}$ such that (i) a large value of the objective in Eq. 1 is obtained (consistent with CC), and (ii) the graph induced by the non-neutral objects is as dense as possible (i.e., most edge weights are $+1$ or $-1$). Any object that hinders either of these goals should be assigned to the neutral set $S_{0}$. This includes ambiguous objects, such as those with significant similarity to multiple clusters or those with inherently weak associations (e.g., low-degree nodes). Importantly, there exists a natural trade-off between the size and density of the non-neutral clusters: small clusters can trivially achieve high density. As we will demonstrate, our objective allows for a flexible balance of this trade-off.
@@
The third term in Eq. 3 has been previously applied to the minimum cut objective for unsigned networks (Chehreghani22_shift, ). In our context (i.e., for the PCD problem), the objective in Eq. 3 achieves two goals simultaneously: it penalizes the formation of (i) large/sparse and (ii) highly imbalanced non-neutral clusters. The second property is easy to see, as for a clustering with $k$ clusters and $n$ objects, the term $\sum_{m\in[k]}|S_{m}|^{2}$ is minimized when each cluster is assigned $\frac{n}{k}$ objects (i.e., the clusters are perfectly balanced). We introduce regularization as an additive term rather than a normalization for two key reasons: (i) It allows a flexible trade-off between the number of non-neutral objects and the density of the graph induced by them (controlled by the parameter $\beta\in\mathbb{R}$), which is a desirable property in this context as discussed in the beginning of this section. This possibility is absent in the existing methods that are based on Eq. 2 (polarity). (ii) It enables the development of an efficient optimization procedure based on local search with strong convergence guarantees. In the context of CC, local search–based approximation algorithms are known to significantly outperform other methods (DBLP:conf/aaai/ThielCD19, ; Chehreghani22_shift, ). In this paper, we develop the first scalable local search algorithm specifically tailored to the PCD setting. In Section 4, we demonstrate across a range of real-world and synthetic datasets that the advantages of local search optimization (e.g., for CC) carry over to PCD as well. We are now ready to formally state our problem, and we subsequently highlight its computational complexity in Thm. 1.
@@ 3.1 Block-Coordinate Frank-Wolfe Optimization
Our objective in Eq. 3 can be written as $\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A_{i,j}$.
We observe that the regularization term in Eq. 3 is equivalent to shifting the intra-cluster similarities by $-\beta$. This reformulation proves highly useful for the remainder of this section. In our context, each object $i\in[n]$ defines a block. We represent the cluster membership of object $i$ using ${\bm{x}}_{i}\in\{{\bm{e}}_{0},\dots,{\bm{e}}_{k}\}$, where ${\bm{e}}_{m}$ (for $m\in\{0,\dots,k\}$) are the standard basis vectors. Each ${\bm{x}}_{i}$ is a vector of dimension $k+1$, with index zero indicating membership in the neutral set $S_{0}$. Specifically, if $x_{i0}=1$, object $i$ is assigned to $S_{0}$. Using this notation, we can now define our objective as follows.
@@ 3.2 Equivalence to a Local Search Approach
4:     Assign object $i$ to a cluster (or make it neutral) which maximally increases our objective in Eq. 3.
@@ 3.4 Improving the Computational Complexity
1:  Randomly assign each object $i\in[n]$ to one of the clusters in $S_{[k]}$, or to the neutral set $S_{0}$
2:  Initialize $X$ based on initial clustering $S_{[k]}$
3:  $M\coloneqq 2AX$
@@
7:     $\phi_{i}\coloneqq\sum_{p}M_{i,p}$
8:     $G_{i,p}\coloneqq-\beta+(1+\alpha)M_{i,p}-2\beta|S_{p}|+2\beta\bm{1}_{[i\in S_{p}]}-\alpha\phi_{i},\ \forall p\in[k]$
@@
11:     If $p^{\ast}=\hat{p}$, then skip to next iteration
@@
In the previous section, we demonstrated that Alg. 2 is guaranteed to converge at the linear rate $O(1/t)$, making it highly efficient. In this section, we propose an alternative version of Alg. 2, designed to enhance the efficiency of each step $t$ while maintaining full equivalence in functionality. This ensures that the convergence analysis from the previous section still remains valid. Firstly, a naive implementation of Alg. 2 has a complexity of $O(Tk^{2}n^{2})$, as each iteration requires $O(k^{2}n^{2})$ to compute the full objective in Eq. 5 for every candidate cluster in order to determine the best cluster for the current object $i$. Since the number of iterations $T$ until convergence is typically larger than $n$, this approach can become computationally expensive. Part (b) of Thm. 2 offers an alternative: instead of evaluating the full objective, we can compute the gradient $G_{i,:}$, which involves only terms related to object $i$. Let $M_{i,m}\coloneqq 2\sum_{j\in S_{m}}A_{i,j}$, $\phi_{i}\coloneqq\sum_{m\in[k]}M_{i,m}$ and $\beta_{im}\coloneqq 2\beta|S_{m}|-2\beta\bm{1}_{[i\in S_{m}]}$. Given this, we present the following theorem.
Theorem 4.
$G_{i,m}=-\beta+(1+\alpha)M_{i,m}-\beta_{im}-\alpha\phi_{i}$ (8)
for all $m\in[k]$ and $G_{i,0}=0$.
A naive calculation of the full gradient $G_{i,:}$ for block $i$ is $O(k^{2}n)$. However, the particular form of the gradient presented in Eq. 8 makes it $O(kn)$. See the proof of Thm. 4 for further insight on this. From Thm. 2, the gradient $G_{i,m}$ represents the impact on the full objective in Eq. 5 if object $i$ is placed in cluster $m$. Thus, because $G_{i,0}=0$, we observe that an object $i$ is made neutral if its contribution to all non-neutral clusters is currently negative. Moreover, the total complexity is now reduced to $O(Tkn)$, which is a significant improvement over the naive approach with complexity of $O(Tk^{2}n^{2})$.
We present a third approach, shown in Alg. 3. We define a matrix $X\in\{0,1\}^{n\times k}$, with $X_{i,m}=1$ if object $i$ belongs to cluster $m\in[k]$, and zero otherwise. Neutral objects $i\in S_{0}$ have rows $X_{i,:}$ of zeros. The procedure precomputes the matrix $M=2AX$, where $M_{i,m}$ is the total similarity of object $i$ to cluster $m$. Precomputing $M$ takes $O(kn^{2})$, but allows gradient computation in $O(k)$ (line 8). We then have to update $M_{i,m}$ accordingly (lines 13 and 14), which is $O(n)$, reducing the per-iteration complexity to $O(n+k)$. Thus, the total complexity is $O(kn^{2}+T(n+k))$, which improves on the $O(Tnk)$ approach because, (i) computing $M$ involves a sparse matrix product, which is highly efficient in practice, and (ii) since $T>n$, reducing per-iteration cost leads to significant practical gains. On the largest datasets in our experiments, Alg. 3 completes in seconds or minutes, while the naive version would take hours or days. A runtime analysis of all three approaches is provided in Appendix G.5.

### v3

@@ Abstract
Signed networks, where edges are labeled as positive or negative to represent friendly or antagonistic interactions, provide a natural framework for analyzing polarization, trust, and conflict in social systems. Detecting meaningful group structures in such networks is crucial for understanding online discourse, political divisions, and trust dynamics. A key challenge is to identify communities that are internally cohesive and externally antagonistic, while allowing for neutral or unaligned vertices. In this paper, we propose a method for identifying $k$ polarized communities that addresses a major limitation of prior methods: their tendency to produce highly size-imbalanced solutions. We introduce a novel optimization objective that avoids such imbalance. In addition, it is well known that approximation algorithms based on local search are highly effective for clustering signed networks when neutral vertices are not allowed. We build on this idea and design the first local search algorithm that extends to the setting with neutral vertices while scaling to large networks. By connecting our approach to block-coordinate Frank-Wolfe optimization, we prove a linear convergence rate, enabled by the structure of our objective. Experiments on real-world and synthetic datasets demonstrate that our method consistently outperforms state-of-the-art baselines in solution quality, while remaining competitive in computational efficiency.
@@ 1 Introduction
The main contributions of this paper are as follows:
  1. (i)
    We propose a novel formulation for the PCD problem that encourages more balanced communities, addressing a key limitation of previous work that typically optimize polarity (DBLP:conf/nips/TzengOG20, ) (see Eq. 2). As demonstrated in our experiments, optimizing polarity often leads to clustering solutions with multiple empty clusters. The importance of promoting balanced communities (to avoid trivial solutions where all objects are placed in one or few clusters) is well established in the graph clustering literature DBLP:conf/cikm/ChiangWD12 . In Appendix E, we expand on this, and provide some examples of practical scenarios where (reasonably) balanced communities are favorable in our context. We note that DBLP:journals/ml/GulloMT24 also proposes an objective for PCD called $\gamma$-polarity aimed at addressing cluster imbalance; however, it is restricted to the case of $k=2$ clusters. In contrast, our objective supports an arbitrary number of clusters $k$. Nonetheless, we compare to their method experimentally for $k=2$, and explain how it differs conceptually in Appendix F.
  2. (ii)
    Motivated by the effectiveness of local search-based approximation algorithms for CC (and many other machine learning models), we propose the first scalable local search algorithm for PCD, which explicitly allows for neutral objects.
  3. (iii)
    We establish a linear convergence rate of our local search algorithm by connecting it to block-coordinate Frank-Wolfe optimization (fw, ; DBLP:conf/icml/Lacoste-JulienJSP13, ). This connection is made possible utilizing the specific structure of our proposed optimization objective and extending the analysis in (DBLP:conf/aaai/ThielCD19, ).
  4. (iv)
    We propose techniques that allow the local search method to scale to large networks.
  5. (v)
    Finally, through extensive experiments on commonly used real-world and synthetic datasets in previous work on PCD, we show that our approach consistently outperforms state-of-the-art baselines in terms of (a) recovering ground-truth solutions and (b) finding high quality solutions of reasonable cluster size balance.
@@ 2.1 Correlation Clustering
We begin by noting that for CC, unlike PCD to be discussed in the next subsection, a clustering $S_{[k]}$ is a partition of $V$, meaning $V=\bigcup_{m\in[k]}S_{m}$ and each $S_{m}$ is disjoint. A notable feature of CC is its ability to automatically determine the number of clusters DBLP:conf/kdd/BonchiGL14 , but here we we focus on the $k$-constrained variant of CC (DBLP:journals/toc/GiotisG06, ) as it is most relevant to our problem. The $k$-$C$ problem can be defined as shown below.
Problem 1 ($k$-$C$).
@@
In other words, the goal is to find a clustering that (i) maximizes intra-cluster similarities and (ii) minimizes inter-cluster similarities. In the CC literature, it is known that maximizing certain subsets of terms in Eq. 1, such as the total number of agreements $N^{+}_{\text{intra}}+N^{-}_{\text{inter}}$, is equivalent to maximizing the full objective ethz-a-010077098 . However, this equivalence does not hold for PCD when neutral objects are allowed, as each of the four terms in Eq. 1 contributes uniquely to the decision of whether an object should be clustered or left neutral. We formally show this in Appendix A, and thus focus on the full objective in the next section when we introduce PCD. CC is known to be NP-hard DBLP:journals/ml/BansalBC04 , and many approximation algorithms have been proposed (DBLP:journals/ml/BansalBC04, ; CharikarGW05, ; DemaineEFI06, ; AilonCN08, ), with local search methods standing out for their strong performance in both clustering quality and computational efficiency (DBLP:conf/aaai/ThielCD19, ; Chehreghani22_shift, ; DBLP:journals/tmlr/AronssonC24, ).
@@ 2.2 Polarized Community Discovery
For PCD, we introduce a neutral set $S_{0}$. As a result, each object in $V$ is either assigned to one of the non-neutral clusters $S_{1},\dots,S_{k}$ or designated as neutral by placing it in $S_{0}$. Consequently, a clustering $S_{[k]}$ is no longer a partition of $V$ and we have $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ (although all clusters are still disjoint). Given this, the goal of PCD is to identify non-neutral clusters $S_{[k]}$ such that (i) a large value of the objective in Eq. 1 is obtained (consistent with CC), and (ii) the graph induced by the non-neutral objects is as dense as possible (i.e., most edge weights are +1+1 or 1-1). Any object that hinders either of these goals should be assigned to the neutral set $S_{0}$. This includes ambiguous objects, such as those with significant similarity to multiple clusters or those with inherently weak associations (e.g., low-degree nodes). Importantly, there exists a natural trade-off between the size and density of the non-neutral clusters: small clusters can trivially achieve high density. As we will demonstrate, our objective allows for a flexible balance of this trade-off.
@@
The third term in Eq. 3 has been previously applied to the minimum cut objective for unsigned networks (Chehreghani22_shift, ). In our context (i.e., for the PCD problem), the objective in Eq. 3 achieves two goals simultaneously: it penalizes the formation of (i) large/sparse and (ii) highly imbalanced non-neutral clusters. The second property is easy to see, as for a clustering with $k$ clusters and $n$ objects, the term $\sum_{m\in[k]}|S_{m}|^{2}$ is minimized when each cluster is assigned $\frac{n}{k}$ objects (i.e., the clusters are perfectly balanced). Notably, the squaring of cluster sizes is what encourages cluster size balance.
We introduce regularization as an additive term rather than a normalization for two key reasons: (i) It allows a flexible trade-off between the number of non-neutral objects and the density of the graph induced by them (controlled by the parameter $\beta\in\mathbb{R}$), which is a desirable property in this context as discussed in the beginning of this section. This possibility is absent in the existing methods that are based on Eq. 2 (polarity). (ii) It enables the development of an efficient optimization procedure based on local search with strong convergence guarantees. In the context of CC, local search–based approximation algorithms are known to significantly outperform other methods (DBLP:conf/aaai/ThielCD19, ; Chehreghani22_shift, ).
In this paper, we develop the first scalable local search algorithm specifically tailored to the PCD setting. In Section 4, we demonstrate across a range of real-world and synthetic datasets that the advantages of local search optimization (e.g., for CC) carry over to PCD as well. We are now ready to formally state our problem, and we subsequently highlight its computational complexity in Thm. 1.
@@ 3.1 Block-Coordinate Frank-Wolfe Optimization
Our objective in Eq. 3 can be written as
$\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A_{i,j}.$ (5)
We observe that the regularization term in Eq. 3 is equivalent to shifting the intra-cluster similarities by $\beta$. This reformulation proves highly useful for the remainder of this section. In our context, each object $i\in[n]$ defines a block. We represent the cluster membership of object $i$ using ${\bm{x}}_{i}\in\{{\bm{e}}_{0},\dots,{\bm{e}}_{k}\}$, where ${\bm{e}}_{m}$ (for $m\in\{0,\dots,k\}$) are the standard basis vectors. Each ${\bm{x}}_{i}$ is a vector of dimension $k+1$, with index zero indicating membership in the neutral set $S_{0}$. Specifically, if $x_{i0}=1$, object $i$ is assigned to $S_{0}$. Using this notation, we can now define our objective as follows.
@@ 3.2 Equivalence to a Local Search Approach
4:  Assign object $i$ to a cluster in $S_{[k]}$, or to the neutral set $S_{0}$, which maximally increases our objective in Eq. 3
@@ 3.4 Improving the Computational Complexity
In the previous section, we demonstrated that Alg. 2 is guaranteed to converge at the linear rate $O(1/t)$, making it highly efficient. In this section, we propose an alternative version of Alg. 2, designed to enhance the efficiency of each step $t$ while maintaining full equivalence in functionality. This ensures that the convergence analysis from the previous section still remains valid. Firstly, a naive implementation of Alg. 2 has a complexity of $O(Tk^{2}n^{2})$, as each iteration requires $O(k^{2}n^{2})$ to compute the full objective in Eq. 6 for every candidate cluster in order to determine the best cluster for the current object $i$. Since the number of iterations $T$ until convergence is typically larger than $n$, this approach can become computationally expensive. Part (b) of Thm. 2 offers an alternative: instead of evaluating the full objective, we can compute the gradient $G_{i,:}$, which involves only terms related to object $i$. Based on this, we present the following theorem.
Theorem 4.
$G_{i,m}=2(1+\alpha)\sum_{j\in S_{m}}A_{i,j}-2\alpha\sum_{p\in[k]}\sum_{j\in S_{p}}A_{i,j}-2\beta|S_{m}|+2\beta\bm{1}[i\in S_{m}]-\beta$ (9)
for all $m\in[k]$ and $G_{i,0}=0$.
A naive calculation of the full gradient $G_{i,:}$ for block $i$ is $O(k^{2}n)$. However, the specific structure of the gradient in Eq. 9 reduces the complexity to $O(kn)$, since the term $\sum_{p\in[k]}\sum_{j\in S_{p}}A_{i,j}$ is independent of the cluster $m$ and can therefore be precomputed (see Alg. 3). See the proof of Thm. 4 for further insight on this. From Thm. 2, the gradient $G_{i,m}$ represents the impact on the full objective in Eq. 6 if object $i$ is placed in cluster $m$. Thus, because $G_{i,0}=0$, we observe that an object $i$ is made neutral if its contribution to all non-neutral clusters is currently negative. Moreover, the total complexity is now reduced to $O(Tkn)$, which is a significant improvement over the naive approach with complexity of $O(Tk^{2}n^{2})$.
@@
1: Randomly assign each object $i\in[n]$ to one of the clusters in $S_{[k]}$, or to the neutral set $S_{0}$.
2: Initialize $X\in\{0,1\}^{n\times k}$, with $X_{i,m}=1$ if object $i$ belongs to cluster $m\in[k]$, and $X_{i,m}=0$ otherwise. Neutral objects $i\in S_{0}$ have rows $X_{i,:}$ of zeros.
[…]
@@
7:  $M_{i}\coloneqq\sum_{p}M_{i,p}$
8:  $G_{i,p}\coloneqq(1+\alpha)M_{i,p}-\alpha M_{i}-2\beta|S_{p}|+2\beta\bm{1}[i\in S_{p}]-\beta,\ \forall p\in[k]$ {See Eq. 9}
@@
11:  if $p^{\ast}=\hat{p}$ then skip to next iteration
@@
We present a third approach, shown in Alg. 3. We define a matrix $X\in\{0,1\}^{n\times k}$, with $X_{i,m}=1$ if object $i$ belongs to cluster $m\in[k]$, and zero otherwise. Neutral objects $i\in S_{0}$ have rows $X_{i,:}$ of zeros. The procedure precomputes the matrix $M=2AX$, where $M_{i,m}$ is the total similarity of object $i$ to cluster $m$. Precomputing $M$ is $O(kn^{2})$, but allows gradient computation in $O(k)$ (line 8). We then have to update $M$ accordingly (lines 13 and 14), which is $O(n)$, reducing the per-iteration complexity to $O(n+k)$. The total complexity is $O(kn^{2}+T(n+k))$, which improves on the $O(Tnk)$ approach because, (i) computing $M$ involves a sparse matrix product, which is highly efficient in practice, and (ii) since $T>n$, reducing per-iteration cost leads to significant practical gains.
On the largest datasets in our experiments, Alg. 3 completes in seconds or minutes, while the naive version would take hours or days. Figure 1 presents a runtime comparison of the three approaches discussed above. The method in Alg. 3 (LSPCD) achieves the best computational efficiency, significantly outperforming the naive approach in Alg. 2. Consequently, Alg. 3 is used in all subsequent experiments. In Appendix G.5, we further demonstrate the scalability of Alg. 3 to large-scale graphs.
Figure 1: Comparison of runtime for the three implementations of Alg. 2 (local search) introduced in Section 3.4 by varying the graph size $n$ and the number of non-neutral clusters $k$, using data generated from the m-SSBM model. See Section 4 for a description of this dataset. The noise level is fixed at $\eta=0.4$. When varying $n$, we fix $k=4$; when varying $k$, we fix $n=5000$. LSPCD corresponds to Alg. 3 and is used in all subsequent experiments because of its superior computational efficiency.